# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With Affine Transformation Contrastive Learning

### v3

@@ 4.4 Downstream Fine-tuning
TABLE I: Semantic Segmentation: Comp. on UAVID
@@
[…]
5 Experiments
TABLE II: Semantic Segmentation: Comp. on UDD6
@@ 5 Experiments
TABLE III: Semantic Segmentation: Comp. on Floodnet
@@
5.1 Pre-training Implementation
In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized. Before pre-training, we initialize the model parameters using Kaiming initialization. Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training. During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images. The mask size is set to 32 pixels. The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs. Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs. The input image size remains 224, and similar random cropping and scaling are applied. The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
5.2 Performance on Single/Multi-modal Downstream Tasks
We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios. Detailed descriptions of the datasets, experimental setups, and additional visualizations for each task are provided in the Appendix.
5.2.1 Semantic Segmantation
Comparative results on the UAVID, UDD6, and FloodNet datasets are presented in Tabs. I, II, and III, respectively. The combination of RingMo-Aerial with Mask2Former achieves segmentation outcomes that significantly outperform those of other methods.
Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks. RingMo-Aerial consistently produces superior segmentation maps. Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures. In particular, the RingMo-Aerial + UperNet + ARS-Adapter and RingMo-Aerial + Mask2Former methods demonstrate higher accuracy and clarity in the segmentation of small targets (e.g., stationary and moving vehicles), with enhanced boundary and detail capture. This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
5.2.2 Object Detection
TABLE IV: Object Detection: Comp. on VisDrone-DET
@@ 5.2.2 Object Detection
TABLE V: Object Detection: Comp. on UAVDT
@@
TABLE VI: Object Detection: Comp. on ShipDataset
@@
TABLE VII: Object Detection: Comp. on DroneVehicle
@@
RingMo-Aerial(C) RGB+IR RingMo-Aerial 71.6
TABLE VIII: Object Detection: Comp. on AerialSARData
@@
An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance. On unimodal datasets such as VisDrone-DET (as shown in Tab.IV) and UAVDT (as shown in Tab.V), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods. The experimental results further demonstrate that RingMo-Aerial, compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
@@
What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.VI), and PP-PicoDet-L follows. In contrast, YOLOX-L and YOLOX-Tiny show the worst results. Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
@@
For the multimodal dataset, DroneVehicle (as shown in Tab.VII), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods. Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial. This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications. Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
@@
In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.VIII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance. Mainly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and a mAP50 of 84.4%, highlighting its superiority in SAR modality. These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery. Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments. These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
@@
Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields. For optical images, as shown in the appendix, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets. In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups. Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms. RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images. As shown in the appendix, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence. The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images. RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for ARS.
5.2.3 Scene Classification
TABLE IX: Scene Classification: Comp. on AID and RESISC-45

### v4

@@ 4.4 Downstream Fine-tuning
$\text{Adapter-MSA}=\text{Up}(\text{GELU}(\text{Down}(x)))$ (19a)
$\text{Adapter-FFN}=\text{Up}(\text{GELU}(\text{Conv}(\text{GELU}(\text{Down}(x)))))$ (19b)
[…]
5 Experiments
5.1 Pre-training Implementation
In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized. Before pre-training, we initialize the model parameters using Kaiming initialization. Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training. During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images. The mask size is set to 32 pixels. The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs. We adopt a masking ratio of 0.6 as the default configuration. This value is selected based on preliminary experiments, balancing semantic abstraction and structural preservation in ARS images characterized by multi-scale sparsity. Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs. The input image size remains 224, and similar random cropping and scaling are applied. The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
Figure 13: Performance improvement of RingMo-Aerial over previous SOTA methods across various remote sensing datasets. The horizontal axis denotes the percentage change in performance.
5.2 Performance on Single/Multi-modal Downstream Tasks
We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios. Fig. 13 presents a comprehensive comparison between RingMo-Aerial and previous state-of-the-art (SOTA) methods on 17 datasets. It clearly demonstrates that RingMo-Aerial achieves superior performance, particularly in ARS imagery characterized by diverse viewing angles and object scales. Notably, RingMo-Aerial yields significant gains in semantic segmentation (e.g., +4.66 mIoU on FloodNet), object detection (e.g., +3.6 mAP on VisDrone-DET), and change detection tasks (e.g., +4.02 F1 on CDD), highlighting its strong generalization and representation capabilities across complex airborne environments.
Detailed descriptions of the datasets, experimental setups, table of comparison with state-of-the-art methods and additional visualizations for each task are provided in the Appendix (Semantic Segmantation C, Object Detection D, Change Detection E, Scene Classification F, 3D-Reconstruction G and Object Tracking H).
5.3 Ablation Experiments
5.3.1 Ablation Experiments on pre-training
TABLE I: Ablation Experiments of pretrain methods
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo 71.6 61.9 28.4 45.2
RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
TABLE II: Ablation Experiments with the proposed ARS-Adapter
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET. The comparative analysis involves three foundation models: RingMo pre-trained with MIM [1], RingMo-Aerial solely pre-trained with MIM, and RingMo-Aerial pre-trained with both MIM and CL. This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
The comparison is shown in Fig.14 and Tab.I. The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models. Notably, RingMo-Aerial trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance. This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter. Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL. Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.II. The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
TABLE III: Comp. between fine-tuning methods
Methods Parameters UAVID VisDrone-DET
mAcc mIoU mAP mAP50
Froezn 0% 61.2 49.3 21.5 36.2
Full 100% 76.5 66.4 31.1 51.8
Adapter 0.39% 68.9 57.7 28.9 49.1
AdaptFormer 0.39% 71.3 62.1 30.3 50.8
ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
The experimental results are presented in Tab.III. For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer). Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment. ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency. Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
5.3.3 Ablation of Mask ratio
TABLE IV: Object detection performance of RingMo-Aerial under different masking ratios on the DroneVehicle dataset (modality: RGB+IR). The result at 0.6 is from the original configuration.
Mask Ratio mAP mAP50 mAP75 mAPs mAPm mAPl
0.5 0.489 0.728 0.566 0.196 0.493 0.664
0.6 0.492 0.735 0.570 0.228 0.497 0.684
0.7 0.487 0.726 0.567 0.195 0.490 0.668
0.8 0.490 0.731 0.567 0.235 0.494 0.681
TABLE V: Semantic segmentation performance of RingMo-Aerial under different masking ratios on the UAVid dataset. The result at 0.6 is from the original configuration.
Mask Ratio mIoU (%) mAcc (%)
0.5 68.29 78.18
0.6 68.70 78.70
0.7 67.72 77.60
0.8 68.56 78.28
To investigate the influence of masking ratio in pretraining, we conduct additional experiments under different ratios (0.5, 0.6, 0.7, 0.8). Results on both object detection (DroneVehicle, RGB+IR) and semantic segmentation (UAVid) are presented in Tab. IV and Tab. V The masking ratio of 0.6 achieves the highest mAP and mIoU, confirming its effectiveness across tasks. Interestingly, higher ratios (e.g., 0.8) slightly improve small object detection, while lower ratios (e.g., 0.5) provide marginal gains on medium-sized targets. These findings suggest that ARS imagery, due to its multi-scale structure and sparse layout, requires more carefully tuned masking strategies than natural images.
5.3.4 Ablation of depthwise convolution placement
TABLE VI: Ablation study on the placement of depthwise convolution in the FE-MSA module. Results are reported on the DroneVehicle dataset with RGB+IR modality.
Variant mAP mAP50 mAP75 mAPs mAPm mAPl
Original* 0.492 0.735 0.570 0.228e 0.497 0.684
Conv1* 0.485 0.720 0.561 0.183 0.485 0.664
Conv2* 0.483 0.719 0.556 0.193 0.485 0.670
    Original: Depthwise convolution is inserted after norm1 and before the attention module (used in our final model). Conv1: Depthwise convolution is placed after the FFN block. Conv2: Depthwise convolution is placed between the multi-head attention and FFN.
To analyze the effect of depthwise convolution placement in the FE-MSA module, we compare three configurations using object detection performance on the DroneVehicle dataset (RGB+IR). As shown in Tab. VI, the results demonstrate that the Original configuration achieves the best overall performance, with 0.492 mAP and consistent advantages across all scales, especially for large objects. Although Conv2 slightly improves small object detection (mAPs = 0.193), it performs worse overall. This confirms that applying depthwise convolution early, before semantic modeling via attention and FFN, provides better spatial enhancement for ARS images. Therefore, we adopt the Original configuration in our final design.
5.3.5 Ablation of depthwise convolution kernel size
TABLE VII: Object detection performance of RingMo-Aerial under different size depthwise separation convolution kernels on the DroneVehicle dataset (modality: RGB+IR). The results at 7 are from the original configuration.
Kernel Size mAP mAP50 mAP75 mAPs mAPm mAPl
3 0.485 0.722 0.564 0.192 0.489 0.661
5 0.490 0.729 0.568 0.179 0.494 0.673
7 0.492 0.735 0.570 0.228 0.497 0.684
9 0.490 0.733 0.566 0.192 0.493 0.684
TABLE VIII: Semantic segmentation performance of RingMo-Aerial under different size depthwise separation convolution kernels on the UAVid dataset. The result at 7 is from the original configuration.
Kernel Size mIoU (%) mAcc (%)
3 67.18 77.77
5 67.57 78.21
7 68.70 78.70
9 67.03 78.15
To investigate the impact of different sizes of depthwise separable convolution kernels during pre-training, we conducted additional experiments using kernels of sizes 3, 5, 7, and 9. The experimental results for object detection (DroneVehicle, RGB+IR) and semantic segmentation (UAVid) are presented in Table. VII and Table. VIII, respectively. The results show that a kernel size of 7 achieves the highest mAP and mIoU, demonstrating its effectiveness across different tasks. These findings suggest that kernels of sizes 3 and 5 provide insufficient receptive fields, limiting the ability to extract complete and informative features. Conversely, a kernel size of 9 results in an excessively large receptive field, introducing redundant information that interferes with the extraction of relevant features. Therefore, a kernel size of 7 strikes the optimal balance, enabling effective feature extraction and leading to superior performance.
6 Conclusion
RingMo-Aerial presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks. Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein. Our experiments demonstrate that RingMo-Aerial achieves SOTA performance across various datasets for ARS vision tasks. This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks. In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research. Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment. Additionally, we will explore the extension of RingMo-Aerial to other vision tasks and scenarios, aiming to expand its application scope.
Acknowledgments
Appendix A Frequency-Enhanced Multi-Head Self-Attention
The following content provides a derivation analysis of the FE-MSA module from both the spatial domain and frequency domain perspectives.
A.1 Spatial-Domain Analysis
Self-attention performs “global” modeling by assigning correlation weights at the patch level. For small targets, fine-scale features remain under-integrated. Inserting a $7\times 7$ convolution before attention effectively enlarges the receptive field at each position from $1\times 1$ (a single point within a patch) to $7\times 7$, while a residual connection preserves the original features:
$Y_{\mathrm{conv}}=X+\mathrm{DWConv}_{7\times 7}(X).$ (20)
For each spatial location $i$,
$y_{i}=x_{i}+\bigl{[}\mathrm{DWConv}_{7\times 7}(X)\bigr{]}_{i},$ (21)
where $\bigl{[}\mathrm{DWConv}_{7\times 7}(X)\bigr{]}_{i}$ aggregates high-frequency (edge, texture) information from a $7\times 7$ neighborhood around $i$, thereby enhancing local edges and structural details of small targets. The convolutional branch thus accentuates small-target features in the local context, enabling subsequent attention to focus more precisely.
Consider the standard self-attention formulation for head $h$:
$A^{(h)}=\mathrm{Softmax}\Bigl{(}Q^{(h)}{K^{(h)}}^{\top}/\sqrt{d}\Bigr{)},$ (22)
$Q^{(h)}=XW_{Q}^{(h)},$ (23)
$K^{(h)}=XW_{K}^{(h)}.$ (24)
After adding the convolutional branch,
$Q^{\prime(h)}$ $=YW_{Q}^{(h)}=XW_{Q}^{(h)}+\mathrm{DWConv}(X)W_{Q}^{(h)},$ (25)
$K^{\prime(h)}$ $=YW_{K}^{(h)}=XW_{K}^{(h)}+\mathrm{DWConv}(X)W_{K}^{(h)}.$ (26)
For any query–key pair $(i,j)$, the modified dot-product logit expands to
$\ell^{\prime}_{ij}$ $=Q^{\prime}_{i}\cdot K^{\prime}_{j}$
$=\underbrace{x_{i}W_{Q}W_{K}^{\top}x_{j}}_{\ell_{ij}}$
$\quad+\underbrace{x_{i}W_{Q}W_{K}^{\top}[\mathrm{Conv}(X)]_{j}+[\mathrm{Conv}(X)]_{i}W_{Q}W_{K}^{\top}x_{j}}_{\Delta_{ij}}$
$\quad+\underbrace{[\mathrm{Conv}(X)]_{i}W_{Q}W_{K}^{\top}[\mathrm{Conv}(X)]_{j}}_{\delta_{ij}}.$ (27)
Here, the original term $\ell_{ij}$ captures pure attention on the raw features, while the enhancement term $\Delta_{ij}$ and the double-enhancement term $\delta_{ij}$ arise from the convolutional branch. Since $\mathrm{Conv}(X)]$ is amplified at high-frequency regions corresponding to small-target edges and textures, these additional terms preferentially boost logits for positions inside or related to small targets. Consequently,
$\ell^{\prime}_{ij}>\ell_{ij}$ (28)
for small-target regions, and the softmax scaling further amplifies attention weights at these locations, resulting in more comprehensive aggregation of small-target information.
A.2 Frequency-Domain Analysis
Denote the 2D Fourier transform of $X\in\mathbb{R}^{H\times W\times C}$ as
$\widehat{X}(\omega_{x},\omega_{y})=\sum_{m=0}^{H-1}\sum_{n=0}^{W-1}X[m,n]e^{-j(\omega_{x}m+\omega_{y}n)}.$ (29)
Small targets concentrate energy at higher spatial frequencies; letting
$E(\omega)$ $=\|\widehat{X}(\omega)\|^{2},$ (30)
$E_{\mathrm{high}}$ $=\int_{\|\omega\|\geq\omega_{0}}E(\omega)\,d\omega$
$>E_{\mathrm{low}}=\int_{\|\omega\|<\omega_{0}}E(\omega)\,d\omega.$ (31)
Patch-level self-attention acts as a low-pass filter, responding strongly to low frequencies while attenuating high frequencies. A rough model is
$H_{\mathrm{attn}}(\omega)=e^{-\alpha\|\omega\|},\quad\alpha>0,$ (32)
or
$H_{\mathrm{att}}(\omega)=\begin{cases}1,&\|\omega\|<\omega_{0},\\ \beta,&\|\omega\|\geq\omega_{0},\quad 0\leq\beta<1.\end{cases}$ (33)
In contrast, a depth-wise separable $7\times 7$ convolution employs a spatial filter $h[m,n]$ whose frequency response
$H_{\mathrm{conv}}(\omega_{x},\omega_{y})=\sum_{m=-3}^{3}\sum_{n=-3}^{3}h[m,n]e^{-j(\omega_{x}m+\omega_{y}n)}$ (34)
acts as a band-pass filter with relatively high gain in the high-frequency range. Therefore, by preceding attention with this convolution, FE-MSA compensates for the attention filter’s high-frequency attenuation:
$\bigl{|}H_{\mathrm{conv}}(\omega)\bigr{|}_{\|\omega\|\geq\omega_{0}}>\bigl{|}H_{\mathrm{attn}}(\omega)\bigr{|}_{\|\omega\|\geq\omega_{0}}.$ (35)
This frequency-domain enhancement better preserves small-target detail during global aggregation.
Appendix B Parameter analysis of ARS-Adapter
Let $C$ be the number of channels (i.e., embedding dimension) of the encoder layer. The Transformer encoder layer includes a depthwise separable convolution layer with $C$ channels, and four linear layers with C channels. The Feed-Forward Network (FFN) layer includes a linear layer with input $C$ and output $4C$, and another linear layer with input $4C$ and output $C$. The total number of parameters in a single encoder layer is then:
$\mathrm{P_{e}}=C\times 7\times 7+4\times C\times C+2\times C\times 4C=49C+12C^{2}$ (36)
Besides, the total parameters added by the ARS-Adapter are calculated as follows:
For each encoder layer, the added components include: Downsampling linear layers $2\times(C\times C/64)$, Upsampling linear layers $2\times(C\times C/64)$ and Depthwise separable convolution layer $\frac{C}{64}\times 3\times 3$. Then, the total additional parameters for each encoder layer are:
$\mathrm{P_{a}}=4\times(C\times C/64)+\frac{C}{64}\times 3\times 3=\frac{9}{64}C+\frac{1}{16}C^{2}$ (37)
Finally, if $C=512$, we can calculate the proportion of added parameters compared to the total parameters of the model:
$\mathrm{Proportion}=\frac{\mathrm{P_{a}}}{\mathrm{P_{e}}+\mathrm{P_{a}}}\approx 0.5\%$ (38)
Appendix C Semantic Segmentation
TABLE IX: Semantic Segmentation: Comp. on UAVID
@@ Appendix C Semantic Segmentation
TABLE X: Semantic Segmentation: Comp. on UDD6
@@
TABLE XI: Semantic Segmentation: Comp. on Floodnet
@@
C.1 Datasets and Evaluation Metrics
C.2 Experiment Settings
C.3 Experiment Results
Comparative results on the UAVID, UDD6, and FloodNet datasets are presented in Tabs. IX,X, and XI, respectively. The combination of RingMo-Aerial with Mask2Former achieves segmentation outcomes that significantly outperform those of other methods.
Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks. As illustrated in Fig.15, RingMo-Aerial consistently produces superior segmentation maps. Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures. In particular, the RingMo-Aerial + UperNet + ARS-Adapter and RingMo-Aerial + Mask2Former methods demonstrate higher accuracy and clarity in the segmentation of small targets (e.g., stationary and moving vehicles), with enhanced boundary and detail capture. This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
Appendix D Object Detection
D.1 Datasets and Evaluation Metrics
TABLE XII: Object Detection: Comp. on VisDrone-DET
@@ D.1 Datasets and Evaluation Metrics
TABLE XIII: Object Detection: Comp. on UAVDT
@@
TABLE XIV: Object Detection: Comp. on ShipDataset
@@
TABLE XV: Object Detection: Comp. on DroneVehicle
@@
RingMo-Aerial(C) RGB+IR RingMo-Aerial 73.5
TABLE XVI: Object Detection: Comp. on AerialSARData
@@ D.3 Experiment Results
An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance. On unimodal datasets such as VisDrone-DET (as shown in Tab.XII) and UAVDT (as shown in Tab.XIII), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods. The experimental results further demonstrate that RingMo-Aerial, compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
@@
What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.XIV), and PP-PicoDet-L follows. In contrast, YOLOX-L and YOLOX-Tiny show the worst results. Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
@@
For the multimodal dataset, DroneVehicle (as shown in Tab.XV), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods. Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial. This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications. Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
@@
In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.XVI), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance. Mainly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and a mAP50 of 84.4%, highlighting its superiority in SAR modality. These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery. Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments. These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
@@
Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields. For optical images, as shown in Fig.16, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets. In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups. Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms. RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images. As shown in Fig.17 and 18, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence. The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images. RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for ARS.
Appendix E Change Detection
E.1 Datasets and Evaluation Metrics
TABLE XVII: Change Detection: Comp. on LEVIR-CD and CDD
Methods LEVIR-CD CDD
F1 IoU F1 IoU
FC-Siam-Di [112] 86.31 83.31 70.61 54.57
FC-Siam-Conc [112] 83.69 76.77 75.11 60.14
IFNet [113] 88.13 82.93 90.30 71.91
STANet [110] 87.30 77.40 84.12 72.22
BIT [114] 89.31 89.37 88.90 80.01
SNUNet [115] 88.16 87.17 83.89 72.11
ChangeFormer [116] 90.40 88.80 89.83 81.53
RingMo [1] 89.53 82.38 90.90 84.09
RingMo-Aerial 92.71 87.10 94.62 90.09
E.2 Experiment Settings
E.3 Experiment Result
The comparison results are shown in Tab.XVII. RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods. In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection. When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in Fig.19. The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images. In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
Appendix F Scene Classification
F.1 Datasets and Evaluation Metrics
TABLE XVIII: Scene Classification: Comp. on AID and RESISC-45